OpenStudy (lncognlto):

Integration question below:

OpenStudy (lncognlto):

\[\int\limits_{ }^{ }\frac{ x }{ (x+1)^2(x^2 +1) }dx\]

OpenStudy (lncognlto):

Now, when I split this question into partial fractions to integrate, I get \[\frac{ A }{ x + 1 } + \frac{ B }{ (x+1)^2 } + \frac{ Cx + D }{ x^2 + 1 }\] while the memo gives \[\frac{ A }{ (x+1)^2 } + \frac{ Bx + C }{ (x^2 + 1) }\]. What rule am I not taking into account?

OpenStudy (fibonaccichick666):

that is weird... I would have split up the x^2+1 too though

OpenStudy (anonymous):

You are doing it right.

OpenStudy (fibonaccichick666):

maybe that modifies somewhere, let me write it out

OpenStudy (anonymous):

First one looks good to me..

OpenStudy (anonymous):

Just find out the constants, you will get the same.

OpenStudy (fibonaccichick666):

Scratch that it's a plus can't use conjugates

zepdrix (zepdrix):

ya your setup looks good Cog :) You have a repeated linear factor, so you have to split up the A and B like that.

OpenStudy (fibonaccichick666):

Just in case you want a secondary reference, but I agree looks good: http://www.purplemath.com/modules/partfrac2.htm

OpenStudy (rsadhvika):

is that a rule when you have square factors in the denominator ?

OpenStudy (fibonaccichick666):

no, the memo is incorrect. You did it by the rules.

OpenStudy (fibonaccichick666):

You may just end up with A=0 though

OpenStudy (anonymous):

@rsadhvika yes, when you got square or cubes, then you do like that.

OpenStudy (rsadhvika):

this looks a lot like \[\large 0.\overline{123} = \frac{1}{10} + \frac{2}{10^2} + \frac{3}{10^3} + \cdots\]

OpenStudy (anonymous):

It is surely the case, if first constant will be \(0\), then first term will vanish, and you are left with second and third only.

OpenStudy (fibonaccichick666):

But, @incognito You should still set it up as you did

OpenStudy (rsadhvika):

i remember reading it somewhere but not really sure how exactly they are related

OpenStudy (fibonaccichick666):

http://www.purplemath.com/modules/partfrac2.htm @rsadhvika , you can read all of the rules here

OpenStudy (anonymous):

I think they are related in your eyes only. :P

OpenStudy (rsadhvika):

it doesn't explain "why" that rule works

OpenStudy (anonymous):

Are partial fractions rule?? Or they are just conventions?

zepdrix (zepdrix):

I don't have a good understanding of why it works myself. There was a really interesting explanation on stackexchange, maybe check it out http://math.stackexchange.com/questions/20963/integration-by-partial-fractions-how-and-why-does-it-work

OpenStudy (rsadhvika):

exactly what i am looking for, thanks @zepdrix !

OpenStudy (fibonaccichick666):

I would say they are a form of factoring. Essentially what you do is break down the bottom into pieces then facilitate the addition which gives you your bottom back, and then you have to solve the system with the original

OpenStudy (lncognlto):

Okay, thanks, I've gotten two of the same coefficients as given in the memo, those for B and C in my expression, but I can't find A and D? all I get is that A + D = 1/2.

OpenStudy (fibonaccichick666):

if you type out the intensive FOILing, I'll check it

OpenStudy (lncognlto):

I rearranged my expression to look like \[\frac{ D }{ x+1 }+\frac{ A }{ (x+1)^2 }+\frac{ Bx + C }{ (x^2 + 1) } = x\] so that the coefficients I am looking for will correspond with the memo, and then I get the expression \[D(x+1)(x^2+1) + A(x^2+1) + (Bx + C)(x+1)^2 = x\] into which I substituted values of x, namely -1, 0, and 1. That's how I've been doing it.

OpenStudy (lncognlto):

From which I got, A = -1/2 and B = 0, and C + D = 1/2.

OpenStudy (anonymous):

@lncognlto please check my method

OpenStudy (fibonaccichick666):

my b which is your A is also -1/2 and My D which is yourC is 1/2, the other two were zero, but instead of inputting x values, you should just solve the system of eqs

OpenStudy (anonymous):

@FibonacciChick666 please check my attachment

OpenStudy (rsadhvika):

thats a clever method!

OpenStudy (anonymous):

That is just a rearrangement, but good use of brain I must say. :)

OpenStudy (anonymous):

It reduces to standard forms.:). we dont ahve to use partial fractions.:)

OpenStudy (fibonaccichick666):

nice rearrangement

OpenStudy (lncognlto):

Okay, I got it. Thanks guys